OpenStudy (anonymous):

A gallon of paint can cover 600 square feet. Determine how many gallons should be purchased to paint two coats on each wall of a room which has dimensions 23 feet by 26 feet and and 7-foot ceiling. Do not include the ceiling – you will only be painting the walls.

OpenStudy (anonymous):

23*26=598 2 gallons

OpenStudy (anonymous):

4 walls: 2 ( 7* 23) + 2 ( 7* 26) = ... 2 coats: 2 * ...

OpenStudy (anonymous):

how do i get the answer

OpenStudy (anonymous):

Can you do it 2 ( 7* 23) + 2 ( 7* 26) = ....?